i am confused about this math question,
two less than a number y is seven times y.
im not sure if the number is y or a different number like x

Answer by nyc_function(2741)

The statement "two less than a number y is seven times y" can be written in the following way: y - 2 = 7y.
You need to solve for y.
y - 2 = 7y
y - 7y = 2
-6y = 2
y = 2/-6
y = -2/6
y = -1/3
Done!